Suppose X~Pois(λ1) and Y~Pois(λ2). Find the conditional mass function for X given X+Y = m
Answer #1

Solution XNP.(41) and Y ~ Pois (12) PIX=> 1 [x+y=mj7 - PIX = 1 n x+y=m) PIX+Y=m) = PIX=71Y=m-r) PIX +Y=m) = P1X=r) Ply=m-r) Pthank you